in a triangle ABC if cos A = 17/22 , cos C =1/14 provve that the ratio of the sides 7:9:11

in a    triangle ABC if cos A = 17/22 , cos C =1/14 provve that the ratio of the sides 7:9:11

Grade:11

2 Answers

Harshit Panwar
6 years ago
cos A=17/22
AB/AC=17/22
 
cos C=1/14
BC/AC=1/14
 
Now we should equate both the denominator.
AB/AC=17×7/22×7
AB/AC=119/154
 
BC/AC=1×11/14×11
BC/AC=11/154
 
THEREFORE,
AB=119 units
BC=11 units
AC=154 units
ratio=119:11:154
 
My answer is 119:11:154
I think that the question is wrong.
Vidyutaa K
5 years ago
Sin a =√195/22
Sin c =√195/14
Sin B = sin(180-A-C)
Sin(A+C)
SinAcosC+cosAsinC
= 9√195/154
a:b:c= 7:9:11